OpenStudy (anonymous):

What is the value of the expression q + 39 when q = –12? A. –51 B. –27 C. 27 D. 51

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

-12+39=39-12=27